1. You need to start a continuous IV infusion of Lasix at 20mg/hour. The concentration of the Lasix infusion is 500mg in 50mL NS. What is the infusion rate in mL/hour?
Answer #1

ANSWER.

   * Calculation of ml/hr.

      ml/hr = Volume(ml) / Time(hr)

         Calculation of volume(ml)

             Volume(ml) = Desired dose / Dose in hand * Quantity

      Here

             Desired dose (D) = 20 mg/hr

             Dose in hand (H) = 500 mg

                     Quantity (Q) = 50 ml

    As per above formula

           Volume(ml) of drug required = D/H * Q

                                                          = 20/500 * 50

                                                          = 0.04 * 50

                                                          = 2 ml.

                                          Time(hr) = 1 hr.

                So

                    Infusion rate (ml/hr) = 2 ml/hr.